Identify the correct dimensions of density and pressure with solutions
1
Expert's answer
2020-02-21T10:06:39-0500

1) Pressure


"\dim P=\frac{\dim F}{\dim A}=\frac{MLT^{-2}}{L^2}=ML^{-1}T^{-2}"

2) Density


"\dim \rho=\frac{\dim m}{\dim V}=\frac{M}{L^3}=ML^{-3}"